A ship of mass 3×10^7kg is at rest it is then pulled to 3 meter by applying a force of 5×10^4 n if friction of water is negligible then find speed of the ship

Grade:9

1 Answers

Arun
6 years ago
 

acceleration of ship a=5*10N/ 3*107 KG

 

                               =5/3 *10-3 m/sec2

 

speed of the ship after movieng 3 m 

 

v2 =u2 +2aS

 

    =0 +2* 5/3 *10-3   * (3)

 

    = 10-2

 

v=.1 m/sec